# Fitness Landscape of Large Language Model-Assisted Automated Algorithm Search

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in algorithm design. However, when integrated into search frameworks for iterative algorithm search, the underlying fitness landscape—critical for understanding search behaviour—remains underexplored. In this paper, we illustrate and analyze the fitness landscape of LLM-assisted Algorithm Search (LAS) using a graph-based approach, where nodes represent algorithms and edges denote transitions between them. We conduct extensive evaluations across six algorithm design tasks and six state-of-the-art LLMs. Our findings reveal that LAS landscapes are highly multimodal and rugged, particularly in combinatorial optimization tasks, with distinct structural variations across tasks and LLMs. For instance, heuristic design tasks exhibit dense clusters of high-performing algorithms, while symbolic regression tasks show sparse, scattered distributions. Additionally, we demonstrate how population size influences exploration-exploitation trade-offs and the evolving trajectory of elite algorithms. These insights not only advance our understanding of LAS landscapes but also provide practical guidance for designing more effective LAS methods.
@@ III Graph-Based Fitness Landscape
In the context of LLM-driven evolutionary search, the search space $S$ comprises candidate algorithms generated by LLMs. The objective function $f$ is defined by specific performance metrics relevant to the task at hand. Since algorithms lack a well-defined distance metric or neighbourhood structure [12], the neighbourhood is instead characterized by correlations between algorithms. Following the approach in [26], we define the neighborhood of an algorithm $a_{i}$ relative to another algorithm $a_{j}$ as follows: $a_{i}\in NS(a_{j})$ if $a_{i}$ is generated from $a_{j}$ (i.e., $a_{i}=LLM(Prompt(a_{j}))$). This formulation enables the introduction of a graph-based representation for illustrating the fitness landscape.
@@ IV-A Date Collection
We tested a diverse set of state-of-the-art LLMs of different sizes and capabilities, including GPT-3.5-Turbo, GPT-4o-Mini, DeepSeek-V3, DeepSeek-Chat, Qwen2-Turbo, and GLM-3-Turbo.
@@ IV-E Landscape using Different Population Sizes
In this section, we present the experimental results of landscape exploration using the DeepSeek-V3 model applied to the TSP task. The visualizations in Figure 6 provide insights into the structural dynamics of the search space as the population size varies. Both 3D and 2D representations are used to illustrate the distribution of candidates (nodes) and their interactions (edges), with edges weighted above 1 highlighted in orange to emphasize significant connections. Overall, smaller population sizes (e.g., 2 and 5) focus on local exploitation due to limited exploration, while larger sizes (e.g., 10 and 50) enable broader exploration, uncovering diverse solutions and revealing the landscape’s inherent complexity with multiple unexplored peaks. The results demonstrate that population size plays a critical role in shaping the search dynamics of LLM-driven evolutionary heuristics.
@@ V Discussion
V-A Landscape Features of LLM-driven Algorithm Search
We identify several key landscape features of the LLM-driven Evolutionary Heuristic Search, which provide insights into its structural and behavioural characteristics:
@@ V-A Landscape Features of LLM-driven Algorithm Search
  • Low Connectivity: The graph representation of the search landscape demonstrates low connectivity, as evidenced by the low average degree and sparsity of connections. This feature suggests that the evolutionary process operates in a fragmented manner, with limited interactions between nodes (solutions). Such low connectivity may impact the efficiency of information exchange and the propagation of beneficial traits across the population.
@@ VI Conclusion
In this paper, we study the fitness landscape of LLM-assisted automated Algorithm Search (LAS) using a graph-based approach. We try to advance our understanding of LLM-assisted algorithm search, offering practical guidance for improving search frameworks and paving the way for more efficient and effective automated algorithm design. Through extensive experiments on six diverse algorithm design tasks and six state-of-the-art LLMs, we uncovered several key insights.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in algorithm design. However, when integrated into search frameworks for iterative algorithm search, the underlying fitness landscape—critical for understanding search behaviour—remains underexplored. In this paper, we illustrate and analyze the fitness landscape of LLM-assisted Algorithm Search (LAS) using a graph-based approach, where nodes represent algorithms and edges denote transitions between them. We conduct extensive evaluations across six algorithm design tasks and six commonly used LLMs. Our findings reveal that LAS landscapes are highly multimodal and rugged, particularly in combinatorial optimization tasks, with distinct structural variations across tasks and LLMs. For instance, heuristic design tasks exhibit dense clusters of high-performing algorithms, while symbolic regression tasks show sparse, scattered distributions. Additionally, we demonstrate how population size influences exploration-exploitation trade-offs and the evolving trajectory of elite algorithms. These insights not only advance our understanding of LAS landscapes but also provide practical guidance for designing more effective LAS methods.
@@ III Graph-Based Fitness Landscape
In the context of LLM-assisted algorithm search, the search space $S$ comprises candidate algorithms generated by LLMs. The objective function $f$ is defined by specific performance metrics relevant to the task at hand. Since algorithms lack a well-defined distance metric or neighbourhood structure [12], the neighbourhood is instead characterized by correlations between algorithms. Following the approach in [26], we define the neighborhood of an algorithm $a_{i}$ relative to another algorithm $a_{j}$ as follows: $a_{i}\in NS(a_{j})$ if $a_{i}$ is generated from $a_{j}$ (i.e., $a_{i}=LLM(Prompt(a_{j}))$). This formulation enables the introduction of a graph-based representation for illustrating the fitness landscape.
@@ IV-A Date Collection
We tested a diverse set of commonly used LLMs of different sizes and capabilities, including GPT-3.5-Turbo, GPT-4o-Mini, DeepSeek-V3, DeepSeek-Chat, Qwen2-Turbo, and GLM-3-Turbo.
@@ IV-E Landscape using Different Population Sizes
In this section, we present the experimental results of landscape exploration using the DeepSeek-V3 model applied to the TSP task. The visualizations in Figure 6 provide insights into the structural dynamics of the search space as the population size varies. Both 3D and 2D representations are used to illustrate the distribution of candidates (nodes) and their interactions (edges), with edges weighted above 1 highlighted in orange to emphasize significant connections. Overall, smaller population sizes (e.g., 2 and 5) focus on local exploitation due to limited exploration, while larger sizes (e.g., 10 and 50) enable broader exploration, uncovering diverse solutions and revealing the landscape’s inherent complexity with multiple unexplored peaks.
@@ V Discussion
V-A Landscape Features of LLM-assisted Algorithm Search
We identify several key landscape features of the LLM-assisted algorithm search:
@@ V-A Landscape Features of LLM-assisted Algorithm Search
  • Low Connectivity: The graph representation of the search landscape demonstrates low connectivity, as evidenced by the low average degree and sparsity of connections. This feature suggests that the search process operates in a fragmented manner, with limited interactions between nodes (solutions). Such low connectivity may impact the efficiency of information exchange and the propagation of beneficial traits across the population.
@@ VI Conclusion
In this paper, we study the fitness landscape of LLM-assisted automated Algorithm Search (LAS) using a graph-based approach. We try to advance our understanding of LLM-assisted algorithm search, offering practical guidance for improving search frameworks and paving the way for more efficient and effective automated algorithm design. Through extensive experiments on six diverse algorithm design tasks and six commonly used LLMs, we uncovered several key insights.